# EXIT: Context-Aware Extractive Compression for Enhancing Retrieval-Augmented Generation

### v2

@@ 1 Introduction
Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Khandelwal et al. (2020) is the task of enhancing Large Language Models (LLMs) responses with relevant external contexts or documents. By grounding answers in evidence, RAG systems have gained much attention for mitigating hallucination issues Ram et al. (2023); Li et al. (2023b) and improving factual reliability Jeong et al. (2024); Xia et al. (2024b).
@@
However, they face significant challenges in practical deployment, as retrieval models sometimes fail to rank the most relevant documents at the top Robertson and Zaragoza (2009); Izacard et al. (2022). One potential solution is to retrieve a larger set of documents to ensure the coverage of the necessary information, but this approach compromises both effectiveness and efficiency. Specifically, for effectiveness, LLMs often struggle with processing long contexts, overlooking critical information located in the middle of contexts Liu et al. (2023). Additionally, irrelevant information in retrieved documents can act as distractors, significantly degrading the overall QA performance Shi et al. (2023a); Li et al. (2023a); Wu et al. (2024). From an efficiency perspective, increasing the context size raises inference latency—due to quadratic complexity in attention computation Xia et al. (2024a)—and API costs tied to input length [^1] https://openai.com/api/pricing/. Moreover, the context window limitations inherent in LLM architectures set strict upper bounds on the maximum input size.
@@
Figure 1: Average QA accuracy (EM) and efficiency (Total Latency) for various compression methods using Contriever-MSMARCO as the retriever and Llama-3.1-8b-Instruct as the reader.
@@
To address these challenges, context compression has emerged as a promising solution, condensing essential information from multiple retrieved contexts through either abstractive or extractive approaches. While they can reduce inference time and filter out irrelevant information, both approaches still have notable drawbacks. Specifically, abstractive compression methods—often implemented via autoregressive generation—summarize or rewrite documents into a single condensed passage Li et al. (2023c); Xu et al. (2024); Yoon et al. (2024); Li et al. (2024); Wang et al. (2023), significantly increasing end-to-end latency due to their token-by-token generation process. For instance, as illustrated in Figure 1, CompAct Yoon et al. (2024), a representative abstractive approach, takes over 8 seconds to process just five documents for a single query, whereas using the original document without compression takes only 1 second.
@@
On the other hand, extractive compression approaches can offer a more efficient alternative Xu et al. (2024); Jiang et al. (2023, 2024), by selecting relevant textual segments (e.g., sentences, or even token-level excerpts) directly from retrieved documents. This strategy reduces both compression time and overall latency. However, current extractive methods have yet to reach their full potential in terms of effectiveness Choi et al. (2021); Pan et al. (2024). They often rely on rigid selection criteria that do not adapt to variations in query complexity or the quality of retrieved documents, and they frequently neglect to fully leverage the broader context when choosing which tokens or sentences to retain. Specifically, as illustrated in Figure 1, while extractive approaches such as RECOMP-Extr Xu et al. (2024) achieve minimal compression time, their inability to dynamically adjust selection processes results in suboptimal QA performance.
@@
Therefore, in this work, we propose a novel compression framework for RAG, EXtractIve ContexT compression (EXIT), designed to enhance both effectiveness and efficiency by improving efficiency through an extractive compression strategy and enhancing effectiveness through dynamic, context-aware sentence selection. Specifically, as shown in Figure 2, EXIT operates in three stages: (1) splitting retrieved documents into sentences, (2) performing parallelizable binary classification (“Yes” or “No”) on each sentence to assess its relevance while considering its full document context, rather than evaluating them independently, and (3) recombining selected sentences while preserving their original order. Therefore, as shown in Figure 1, EXIT frames context compression as a sentence classification problem, enabling it to outperform both compression methods and the uncompressed baseline in terms of speed. Specifically, it reduces processing time from several seconds to about 1 second. Moreover, by leveraging context-aware and adaptive sentence selection, EXIT also surpasses other extractive methods in accuracy. Also, we note that EXIT operates as a plug-and-play module that can be seamlessly integrated into any existing RAG pipeline without architectural modifications.
@@
  • We identify and address the key weaknesses of existing context compression methods: abstractive approaches incur prohibitive latency, while traditional extractive methods rely on rigid, non-adaptive content selection.
@@ 2 Related Work
Retrieval-Augmented Generation.  The RAG pipeline typically follows a naive retrieve-then-generate process, where a single-step retrieval precedes generation Lewis et al. (2020); Shi et al. (2023b); Ram et al. (2023). However, a simple single-step retrieval often fails to rank relevant documents at the top and struggles to handle multi-hop queries requiring multiple pieces of information. To address these, RAG pipelines have evolved into iterative, recursive, and multi-hop retrieval approaches Shao et al. (2023); Trivedi et al. (2023); Khattab et al. (2022), which require multiple retrievals for a single query. While these methods improve information coverage, they also increase end-to-end latency from retrieval to generation, reducing the overall efficiency of the pipeline. Moreover, increasing the document length or the number of retrieved documents as alternative solutions to ensure coverage further exacerbates efficiency issues in these complex retrieval approaches. This not only heightens inference costs Xia et al. (2024a) but also makes it harder to focus on critical details within the documents Liu et al. (2023). In response, reducing the tokens in the retrieved documents while keeping key information in them has gotten attention to complement the efficiency issue of the RAG pipeline.
@@
Context Compression.  Context compression has emerged as a practical remedy for handling increasingly large prompt lengths in RAG pipelines. Existing approaches commonly fall into soft or hard compression. Soft compression focuses on shortening embedding vectors at the token level Wingate et al. (2022); Mu et al. (2023); Ge et al. (2024); Chevalier et al. (2023); Cheng et al. (2024), but requires extensive training and architectural changes, making it unsuitable for black-box LLMs.
@@
Hard compression, by contrast, removes non-essential textual content directly Li et al. (2023d); Jiang et al. (2023), offering a plug-and-play solution compatible even with API-based models such as ChatGPT OpenAI (2023). Hard compression techniques are further divided into abstractive and extractive methods. Abstractive methods employ autoregressive models to generate query-focused summaries, thus drastically reducing token counts at the cost of additional latency and potential hallucinations Zhao et al. (2020). For example, RECOMP-Abst Xu et al. (2024) uses a T5-based summarizer for token reduction but requires dataset-specific training and slows inference. CompAct Yoon et al. (2024) and Refiner Li et al. (2024) take this approach further by leveraging even larger LLMs with 7B parameters, compounding latency issues and increasing resource demands.
Extractive methods select salient segments (e.g., sentences, or tokens) directly from the retrieved documents. This avoids the autoregressive bottleneck and mitigates hallucinations. RECOMP-Extr Xu et al. (2024) is one such example, but its static and context-agnostic selection of only a few sentences per document limits its performance. Similarly, token-level approaches such as LLMLingua family Li et al. (2023d); Jiang et al. (2023); Pan et al. (2024); Jiang et al. (2024) can distort semantic coherence by removing key entities or splitting essential facts.
@@
In short, abstractive methods offer strong compression but suffer from latency and potential halluciantions, while extractive methods are often rigid and lack context awareness. Our work addresses these limitations by proposing a parallelizable, context-aware extractive compression framework. It adaptively selects sentences at scale, preserving semantic integrity and efficiently balancing accuracy with speed, even in complex, multi-step retrieval scenarios.
@@ 3.1 Problem Formulation
Objectives of Compression. For effective and efficient compression in the RAG pipeline, three key criteria should be satisfied: (1) $D^{\prime}$ should contain fewer tokens, as fewer tokens lead to shorter answer generation times (i.e., reading time); (2) $D^{\prime}$ should retain the essential evidence required to answer the query, ensuring effectiveness (i.e., accurate answer); and (3) the compression process should be sufficiently fast enough to avoid significantly increasing the overall end-to-end inference latency.
@@ 3.2 Extractive Context Compression (EXIT)
To achieve three objectives of the compression step, EXIT consists of three main components: sentence-level decomposition, context-aware relevance classification, and document reassembly.
Sentence-Level Decomposition. In Step 1 of Figure 2, EXIT divides each retrieved document into individual sentences using a rule-based sentence tokenizer. For each document $d_{i}\in D$, we produce a sentence set $S_{i}=\{s_{i1},s_{i2},\dots,s_{in}\}$, where $s_{ij}$ is the $j$-th sentence in document $i$. Operating at the sentence level avoids the fragmentation of key phrases and preserves entity relationships that token-level compression techniques Jiang et al. (2023) often disrupt. As a result, the compressed context preserves both syntactic coherence and semantic integrity, ensuring that key information is effectively retained.
Context-Aware Relevance Classification. To effectively and efficiently filter sentences in $D$ that contain key evidence for answering the question, we design two key components for sentence relevance evaluation: document consideration and single-token prediction. First, incorporating the entire document $d_{i}$ is essential, as key evidence may be distributed throughout the context rather than confined to a single sentence, ensuring no critical information is missed and enabling effective context compression. Furthermore, as multiple-token generation proposed in the previous work Yoon et al. (2024); Li et al. (2024) could compromise the overall efficiency of the RAG pipeline, we design the lightweight relevance calculation with only single-token prediction with “Yes” and “No” from the query $q$, document $d_{i}$, and sentence $s_{ij}$. In detail, for each candidate sentence $s_{ij}$, the evaluation model calculates the relevance score $r_{ij}$ of the sentence $s_{ij}$ for a given query $q$ and document $d_{i}$ as follows:
@@
where $P(\cdot|\cdot)$ denotes the likelihood of each token from the evaluation model. This relevance calculation is parallelized across multiple sentences, allowing them to be evaluated simultaneously.
Then, among the sentences in $D$, EXIT selects sentences with a relevance score exceeding a pre-defined threshold $\tau$, as high relevance scores indicate that a sentence contains critical information. Notably, this selection process results in an adaptive number of sentences in the compressed set $D^{\prime}$, rather than a fixed amount. This adaptive approach aligns with prior work Jeong et al. (2024), recognizing that the complexity of queries and the amount of key information vary across queries. As a result, our sentence selection strategy enables effective compression while ensuring all key evidence is included in the compressed set.
Document Reassembly. As shown in Step 3 of Figure 2, EXIT reconstructs the compressed document $D^{\prime}$ by concatenating only the selected sentences in their original order. Following Hwang et al. (2024), preserving the canonical sentence sequence maintains logical flow and contextual coherence. This approach ensures that the resulting compressed document remains understandable and supports accurate downstream reasoning.
@@ 3.3 Classifier Model Training
Training Strategy. Our goal is to train a relevance classifier capable of accurately identifying which sentences provide the evidence required to answer a query. To approximate real-world complexity, we utilize a question-answering dataset that requires multi-sentence reasoning and offers explicit sentence-level annotations of essential information. Leveraging these annotations, we model three typical retrieval outcomes: (1) content that directly provides the needed evidence, (2) passages that appear relevant but lack crucial details, and (3) texts that are entirely irrelevant.
Data Sampling. From the annotated dataset, we draw positive examples from sentences explicitly marked as necessary for producing the correct answer. Within the same documents, we select hard-negative examples—sentences that appear related to the query but do not contain the required evidence—thereby simulating plausible but incomplete retrieval scenarios. Additionally, we sample random negatives from unrelated queries, ensuring the classifier learns to dismiss off-topic content. By maintaining a balanced mix of positives, hard negatives, and random negatives, we create a training set that captures a wide spectrum of retrieval conditions.
@@ 4 Experiment Setups
We conduct comprehensive experiments to evaluate EXIT’s effectiveness and efficiency in context compression for RAG systems. More implementation details of our experiments are in Appendix A.
@@
Baselines.  We compare EXIT against following context compression approaches: 1) Original Documents serves as uncompressed baseline. For abstractive methods, 2) RECOMP-Abs Xu et al. (2024) uses T5-based (775M) summarization tuned for NQ, TQA, HQA (HQA model for 2WIKI), 3) CompAct Yoon et al. (2024) implements Mistral-7B-based iterative compression with 5-segment blocks, and 4) Refiner Li et al. (2024) uses Llama2-7B-based compression. For extractive methods, 5) RECOMP-Extr Xu et al. (2024) employs Contriever-based (110M) sentence-level extraction tuned for NQ, TQA, HQA (HQA model for 2WIKI), and 6) LongLLMLingua Jiang et al. (2024) uses Llama2-7B-chat for token-level extraction with 0.4 dynamic compression rate.
Evaluation Metrics.  We evaluate our approach using three metrics: Exact Match (EM) and F1 score to measure effectiveness in question answering, and end-to-end inference latency (Lat.) in seconds to assess efficiency. Here, end-to-end latency is defined as the total time encompassing both the compression and generation steps, as the retrieval step remains consistent across all methods.
Implementation Details.  We conduct retrieval over the December 2018 Wikipedia dump and apply SpaCy for sentence splitting. We employ vLLM v0.5.5 Kwon et al. (2023) for accelerated inference with the hyperparameter $T=0.0$ and Top-$P=1.0$. We empirically set the relevance threshold $\tau=0.5$. All experiments are conducted on A100-SXM4-80GB GPUs.

### v3

@@ 1 Introduction
Retrieval-Augmented Generation (RAG) Lewis et al. (2020b); Khandelwal et al. (2020) is the task of enhancing the responses of Large Language Models (LLMs) with relevant external contexts or documents. By grounding answers in evidence, RAG systems have gained much attention for mitigating hallucination issues Ram et al. (2023); Li et al. (2023c) and improving factual reliability Jeong et al. (2024); Xia et al. (2024b).
@@
However, they face significant challenges in practical deployment, as retrieval models sometimes fail to rank the most relevant documents at the top Robertson and Zaragoza (2009); Izacard et al. (2022). One potential solution is to retrieve a larger set of documents to ensure the coverage of the necessary information, but this approach compromises both effectiveness and efficiency. Specifically, for effectiveness, LLMs often struggle with processing long contexts, overlooking critical information located in the middle of contexts Liu et al. (2024). Additionally, irrelevant information in retrieved documents can act as distractors, significantly degrading the overall QA performance Shi et al. (2023); Li et al. (2023b); Wu et al. (2024). From an efficiency perspective, increasing the context size raises inference latency—due to quadratic complexity in attention computation Xia et al. (2024a)—and API costs tied to input length [^1] https://openai.com/api/pricing/. Moreover, the context window limitations inherent in LLM architectures set strict upper bounds on the maximum input size Peng et al. (2024).
@@
Figure 1: Average QA accuracy (EM) and efficiency (Total Latency) across various compression methods using Contriever-MSMARCO as the retriever and Llama-3.1-8b-Instruct as the reader. Experiments were conducted on a single A100-80GB GPU, and latency (in seconds) was averaged over all samples from these datasets.
@@
To address these challenges, context compression has emerged as a promising solution, condensing essential information from multiple retrieved contexts through either abstractive or extractive approaches. While they can reduce inference time and filter out irrelevant information, both approaches still have notable drawbacks. Specifically, abstractive compression methods—often implemented via autoregressive generation—summarize or rewrite documents into a single condensed passage Li et al. (2023d); Xu et al. (2024); Yoon et al. (2024); Li et al. (2024); Wang et al. (2023), significantly increasing end-to-end latency due to their token-by-token generation process. For instance, as illustrated in Figure 1, CompAct Yoon et al. (2024), a representative abstractive approach, takes over 8 seconds to process just five documents for a single query, whereas using the original document without compression takes only 1 second.
@@
On the other hand, extractive compression approaches can offer a more efficient alternative Xu et al. (2024); Jiang et al. (2023, 2024), by selecting relevant textual segments (e.g., sentences, or even token-level excerpts) directly from retrieved documents. This strategy reduces both compression time and overall latency. However, current extractive methods have yet to reach their full potential in terms of effectiveness Choi et al. (2021); Pan et al. (2024). They often rely on rigid selection criteria that do not adapt well to variations in query complexity or the quality of retrieved documents, and they frequently neglect to fully leverage the broader context when choosing which tokens or sentences to retain. Specifically, as illustrated in Figure 1, while extractive approaches such as RECOMP-Extr Xu et al. (2024) achieve minimal compression time, their inability to dynamically adjust selection processes results in suboptimal QA performance.
@@
Therefore, in this work, we propose a novel compression framework for RAG, EXtractIve ContexT compression (EXIT), designed to enhance both effectiveness and efficiency by improving efficiency through an extractive compression and by enhancing effectiveness through dynamic, context-aware sentence selection. Specifically, as shown in Figure 2, EXIT operates in three stages: (1) splitting retrieved documents into sentences, (2) performing parallelizable binary classification (“Yes” or “No”) on each sentence to assess its relevance while considering its full document context, rather than evaluating them independently, and (3) recombining selected sentences while preserving their original order. Therefore, as shown in Figure 1, EXIT frames context compression as a sentence classification problem, enabling it to outperform both compression methods and the uncompressed baseline in terms of speed. Specifically, it reduces processing time from several seconds to about 1 second. Moreover, by leveraging context-aware and adaptive sentence selection, EXIT also surpasses other extractive methods in accuracy. Also, we note that EXIT operates as a plug-and-play module that can be seamlessly integrated into any existing RAG pipeline without architectural modifications.
@@
  • We identify and address the key weaknesses of existing context compression methods: abstractive methods incur prohibitive latency, while traditional extractive methods rely on rigid, non-adaptive content selection.
@@ 2 Related Work
Retrieval-Augmented Generation.  RAG enhances LLMs by grounding responses in retrieved external documents Lewis et al. (2020b); Shi et al. (2024); Ram et al. (2023). However, increasing retrieved content often leads to longer contexts, higher inference costs Xia et al. (2024a), and retrieval noise Shi et al. (2023); Liu et al. (2024), impacting efficiency and accuracy.
To mitigate these challenges, ad-hoc methods have been proposed at different stages of the retrieval pipeline, including pre-retrieval and post-retrieval methods. Pre-retrieval methods, such as Landmark Embedding Luo et al. (2024) and Late Chunking Günther et al. (2024), improve document representations to enhance retrieval effectiveness but lack adaptive filtering mechanisms after retrieval and may be difficult to integrate orthogonally with existing retrieval methods. Post-retrieval methods, including reranking Nogueira and Cho (2019); Qin et al. (2023); Li et al. (2023a) and context compression Xu et al. (2024); Yoon et al. (2024); Li et al. (2024); Jiang et al. (2024), refine retrieved documents through reordering or pruning. However, most post-retrieval methods overlook query complexity and operate on a fixed number of retrieved items, limiting their adaptability in balancing effectiveness and efficiency.
We introduce EXIT, a novel post-retrieval compression framework that adaptively filters query-relevant sentences while discarding redundancy. Unlike existing methods, EXIT dynamically adjusts to query complexity and retrieval quality, balancing efficiency and effectiveness. Its lightweight, adaptive design seamlessly integrates into RAG pipelines while remaining orthogonal to pre-retrieval methods.
@@
Context Compression.  Context compression is a practical remedy for handling increasingly extensive prompt lengths in RAG pipelines. Existing methods commonly fall into soft or hard compression. Soft compression focuses on embedding prompts into compact continuous representations Wingate et al. (2022); Mu et al. (2023); Ge et al. (2024); Chevalier et al. (2023); Cheng et al. (2024), but requires extensive training and architectural changes, making it unsuitable for black-box LLMs. Hard compression, by contrast, removes non-essential textual content directly Li et al. (2023e); Jiang et al. (2023), offering a plug-and-play solution even compatible with API-based models such as ChatGPT OpenAI (2023).
@@
Hard compression are further divided into abstractive and extractive methods. Abstractive methods employ autoregressive models to generate query-focused summaries, thus reducing token counts at the cost of additional latency and potential hallucinations Zhao et al. (2020). For example, RECOMP-Abst Xu et al. (2024) uses a T5-based summarizer for token reduction but requires dataset-specific training and slows inference. CompAct Yoon et al. (2024) and Refiner Li et al. (2024) take this method further by leveraging even larger LLMs with 7B parameters, compounding latency issues and increasing resource demands. Extractive methods, such as RECOMP-Extr Xu et al. (2024), directly select salient segments (e.g., sentences or tokens) from retrieved documents. This avoids the autoregressive bottleneck and mitigates hallucinations. However, their static and context-agnostic selection of only a few sentences per document limits its performance. Similarly, token-level methods such as LLMLingua family Li et al. (2023e); Jiang et al. (2023); Pan et al. (2024); Jiang et al. (2024) can distort semantic coherence by removing key entities or splitting essential facts.
@@
In short, abstractive methods offer strong compression but suffer from latency and potential hallucinations, while extractive methods are often rigid and lack context awareness. Our work addresses these limitations by proposing a parallelizable, context-aware extractive compression framework. It adaptively selects sentences at scale, optimizing accuracy and speed in complex retrieval scenarios.
@@ 3.1 Problem Formulation
[…]
While token count influences the reader’s reading time, it alone does not capture the full latency of the pipeline. In practice, end-to-end latency also includes the time spent on the compression step itself. Therefore, an efficient RAG system must minimize both the size of the input and the total time taken from retrieval to answer generation.
@@ 3.2 Extractive Context Compression (EXIT)
To achieve three objectives of the compression step, EXIT consists of three main components: sentence-level decomposition, context-aware relevance classification, and document reassembly. Importantly, EXIT functions as a plug-and-play module that integrates seamlessly into existing RAG pipelines, independent of the specific retriever or reader model used.
Sentence-Level Decomposition. In Step 1 of Figure 2, EXIT divides each retrieved document into individual sentences using a rule-based sentence tokenizer. For each document $d_{i}\in D$, it produces a sentence set $S_{i}=\{s_{i1},s_{i2},\dots,s_{in}\}$, where $s_{ij}$ is the $j$-th sentence in document $i$. Operating at the sentence level avoids the fragmentation of key phrases and preserves entity relationships that token-level compression techniques Jiang et al. (2023) often disrupt. As a result, the compressed context preserves both syntactic coherence and semantic integrity, ensuring that key information is retained.
Context-Aware Relevance Classification. To efficiently filter sentences in $D$ that contain key evidence for answering a query, we introduce a relevance evaluation method based on context awareness and single-token prediction. First, incorporating the entire document $d_{i}$ as context is essential, as understanding a sentence often requires the broader document context rather than an isolated sentence. This ensures that no relevant information is overlooked and enables more effective compression. Second, to maintain efficiency, we adopt lightweight single-token prediction instead of multi-token generation Yoon et al. (2024); Li et al. (2024), which can introduce computational overhead. Inspired by Zhong et al. (2022), we leverage probabilistic classification for sentence relevance assessment. Given query $q$, document $d_{i}$, and sentence $s_{ij}$, the evaluation model predicts relevance using a binary classification with “Yes" and “No" labels:
@@
where $P(\cdot|\cdot)$ represents the likelihood from the evaluation model. This computation is parallelized across sentences for efficiency.
EXIT then selects sentences with a relevance score above a predefined threshold $\tau$, ensuring that only key information is retained. Unlike fixed-size selection, our framework produces an adaptive number of sentences in the compressed set $D^{\prime}$, aligning with prior work Jeong et al. (2024) that acknowledges query-dependent information needs. This approach optimizes compression while preserving essential evidence.
Document Reassembly. As shown in Step 3 of Figure 2, EXIT reconstructs the compressed document $D^{\prime}$ by concatenating selected sentences in their original order, preserving coherence and logical flow Hwang et al. (2024) for accurate downstream reasoning.
@@ 3.3 Classifier Model Training
Training Strategy. Our goal is to train a relevance classifier capable of accurately identifying which sentences provide the evidence required to answer a query. To approximate real-world complexity, we utilize a question-answering dataset that requires multi-sentence reasoning and offers explicit sentence-level annotations of essential information. Leveraging these annotations, we model three typical retrieval outcomes: (1) sentences containing necessary evidence, (2) seemingly relevant sentences missing key details, and (3) entirely irrelevant sentences.
Data Sampling. To train a robust relevance classifier, we construct a diverse dataset reflecting various post-retrieval conditions. Positive samples include sentences essential for correct answers, while hard negatives consist of remaining sentences from the same documents. Additionally, random negatives pair queries with unrelated sentences, helping the model distinguish relevance from noise. This balanced sampling enables effective filtering of non-essential information without relying on explicit supervision.
@@ 4 Experiment Setups
We conduct comprehensive experiments to evaluate EXIT’s effectiveness and efficiency in context compression for RAG systems. More implementation details of our experiments are in Appendix B.
@@
Baselines.  We compare EXIT against the following context compression methods: 1) Original Documents serves as uncompressed baseline. For abstractive methods, 2) RECOMP-Abs Xu et al. (2024) uses T5-based (775M) summarization tuned for NQ, TQA, HQA (HQA model for 2WIKI), 3) CompAct Yoon et al. (2024) implements Mistral-7B-based iterative compression with 5-segment blocks, and 4) Refiner Li et al. (2024) uses Llama2-7B-based compression. For extractive methods, 5) RECOMP-Extr Xu et al. (2024) employs Contriever-based (110M) sentence-level extraction tuned for NQ, TQA, HQA (HQA model for 2WIKI), and 6) LongLLMLingua Jiang et al. (2024) uses Llama2-7B-chat for token-level extraction with 0.4 dynamic compression rate.
Evaluation Metrics.  We use three metrics: Exact Match (EM) and F1 score to measure effectiveness in question answering, and end-to-end inference latency (Lat.) in seconds to assess efficiency. End-to-end latency captures the total wall-clock time required to generate an answer for a single query, including both compression time and LLM inference time. We report the average latency per sample, computed over the entire evaluation set. This provides a more holistic efficiency metric than token count alone, as it directly reflects the practical runtime performance of the full RAG pipeline.
Implementation Details.  We perform document retrieval using the December 2018 Wikipedia dump and apply SpaCy for sentence segmentation. The LLM outputs are generated with a temperature of 0.0 and a top-$P$ of 1.0. We set the relevance threshold $\tau$ to 0.5 based on empirical tuning. All experiments are conducted on A100-SXM4-80GB GPUs.